Suppose Qd=-40*P+963, Qs=37*P+11 for home and Qd*=-40*P+363, Qs*=37*P+35 for foreign. What is the free trade volume to two decimal places?
Answer #1

Od- uop +963 Od* – Yol+363 Os - 370 +11 05-371435 Odds od dst – yo pt 963 =B2+1 YoP+363= 32+35. 952-77 328 = 770 po 952 12.36